# Proactive and Reactive Constraint Programming for Stochastic Project Scheduling with Maximal Time-Lags

### v1

@@ C.1 Feasibility Ratio
Table 10: Feasibility ratio for noise factor $c=1$ We use abbreviations for the proactive approach with $\gamma=0.9$ ($\mathsf{pro}_{0.9}$), and the SAA variant ($\mathsf{pro}_{\mathsf{SAA}}$).
@@
Table 11: Feasibility ratio for noise factor $c=2$ We use abbreviations for the proactive approach with $\gamma=0.9$ ($\mathsf{pro}_{0.9}$), and the SAA variant ($\mathsf{pro}_{\mathsf{SAA}}$).
C.2 Partial Ordering Solution Quality
@@
  • First, we discuss partial orderings for the noise level $c=1$, which are visualized in Figure 6. For $c=1$, the partial ordering shows the same pattern across the different instance sets. However, there are a few exceptions:
@@
  • Then, we analyze the partial orderings for the noise level $c=2$, in Figure 7. Here there are a few things to observe:
    • We observe a shift in the partial orderings for $c=2$.
    • For j10-j20, the partial ordering is similar to $c=1$, although for $c=2$ all arrows are solid (indicating a relation obtained by the Wilcoxon test).
@@
Figure 6: Partial ordering on solution quality for $c=1$. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
Figure 7: Partial ordering on solution quality for $c=2$. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
C.3 Partial Ordering Time Offline
@@
  • Now, we observe the different partial orderings obtained for $c=1$:
@@
  • For noise level $c=2$, we observe in Figure 9:
    • For j10 and j20, the partial ordering is almost the same as for the $c=1$ level, although the flipping already occurs at j20, where the $\mathsf{stnu}$ becomes faster than the $\mathsf{proactive}_{\mathsf{SAA}}$.
@@
Figure 8: Partial ordering time offline c=1. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
Figure 9: Partial ordering time offline c=2. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
C.4 Partial Ordering Time Online
@@
  • For $c=1$ (Figure 11), we find the same partial ordering for each instance set, which is also in line with our expectations:
@@
  • For $c=2$ (Figure 11), we find the same pattern as for $c=1$, for problem sets j10-j30, however, we find that for ubo50 and ubo100, the $\mathsf{stnu}$ starts to outperform other methods, explained by the higher feasibility ratio of the $\mathsf{stnu}$.
@@
Figure 10: Partial ordering time online c=1, j10, j20, j30, ubo50, and ubo100. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
@@
Figure 11: Partial ordering time online c=2, ubo50, and ubo100. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
C.5 Magnitude Tests
@@
In general, we observed that for the pairs of methods for which we found a significant partial ordering, we also found a significant difference in the magnitude of the different metrics. The test results for solution quality are presented in Table 13 (c=1) and 15 (c=2), for time offline in Table 17 (c=1) and 19 (c=2).
The remarkable things that we observed in the results of the magnitude tests, were also already observed while comparing Wilcoxon test results with proportional test results in the earlier analysis for a few cases:
@@
  • For time offline, $c=2$, ubo50 and ubo100, the Wilcoxon test found that the $\mathsf{stnu}$ is better than e.g. $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$, while looking at the magnitude results on double hits (this is an important difference in test procedure!), we find that the $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ are much faster offline than $\mathsf{stnu}$, which is also expected. Then, we observe no difference between $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$, which can be explained by the fact that both have the same offline procedure and we select only double hits for the test.

### v2

@@ E.1 Feasibility Ratio
Table 12: Feasibility ratio for noise factor $\epsilon=1$ We use abbreviations for the proactive approach with $\gamma=0.9$ ($\mathsf{pro}_{0.9}$), and the SAA variant ($\mathsf{pro}_{\mathsf{SAA}}$).
@@
Table 13: Feasibility ratio for noise factor $\epsilon=2$ We use abbreviations for the proactive approach with $\gamma=0.9$ ($\mathsf{pro}_{0.9}$), and the SAA variant ($\mathsf{pro}_{\mathsf{SAA}}$).
E.2 Partial Orderings
E.2.1 Partial Ordering Solution Quality
@@
  • First, we discuss partial orderings for the noise level $\epsilon=1$, which are visualized in Figure 11. For $\epsilon=1$, the partial ordering shows the same pattern across the different instance sets. However, there are a few exceptions:
@@
  • Then, we analyze the partial orderings for the noise level $\epsilon=2$, in Figure 12. Here there are a few things to observe:
    • We observe a shift in the partial orderings for $\epsilon=2$.
    • For j10-j20, the partial ordering is similar to $\epsilon=1$, although for $\epsilon=2$ all arrows are solid (indicating a relation obtained by the Wilcoxon test).
@@
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j10
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j20
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j30
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$ubo50
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$ubo100
Figure 11: Partial ordering on solution quality for $\epsilon=1$. The solid arrow indicates a significant difference obtained by the Wilcoxon test, and the dashed arrow shows only a proportional difference.
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j10
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j20
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j30
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$ubo50
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$ubo100
Figure 12: Partial ordering on solution quality for $\epsilon=2$. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
E.2.2 Partial Ordering Time Offline
@@
  • Now, we observe the different partial orderings obtained for $\epsilon=1$:
@@
  • For noise level $\epsilon=2$, we observe in Figure 14:
    • For j10 and j20, the partial ordering is almost the same as for the $\epsilon=1$ level, although the flipping already occurs at j20, where the $\mathsf{stnu}$ becomes faster than the $\mathsf{proactive}_{\mathsf{SAA}}$.
@@
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j10
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j20
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j30
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$ubo50
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$ubo100
Figure 13: Partial ordering time offline $\epsilon=1$. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j10
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j20
$\mathsf{stnu}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{proactive}_{\mathsf{0.9}}$j30
$\mathsf{stnu}$$\mathsf{proactive}_{\mathsf{0.9}}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$ubo50
$\mathsf{stnu}$$\mathsf{proactive}_{\mathsf{0.9}}$$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$ubo100
Figure 14: Partial ordering time offline $\epsilon=2$. The solid arrow indicates a significant difference obtained by the Wilcoxon test, and the dashed arrow shows only a proportional difference.
E.2.3 Partial Ordering Time Online
@@
  • For $\epsilon=1$ (Figure 16), we find the same partial ordering for each instance set, which is also in line with our expectations:
@@
  • For $\epsilon=2$ (Figure 16), we find the same pattern as for $\epsilon=1$, for problem sets j10-j30, however, we find that for ubo50 and ubo100, the $\mathsf{stnu}$ starts to outperform other methods, explained by the higher feasibility ratio of the $\mathsf{stnu}$.
$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{stnu}$$\mathsf{proactive}_{\mathsf{0.9}}$j10-ubo100
@@
Figure 15: Partial ordering time online $\epsilon=1$, j10, j20, j30, ubo50, and ubo100. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{stnu}$$\mathsf{proactive}_{\mathsf{0.9}}$j10-j30
$\mathsf{reactive}$$\mathsf{proactive}_{\mathsf{SAA}}$$\mathsf{stnu}$$\mathsf{proactive}_{\mathsf{0.9}}$ubo50-ubo100
@@
Figure 16: Partial ordering time online $\epsilon=2$, ubo50, and ubo100. Solid arrow indicating a significant difference obtained by the Wilcoxon test, and dashed arrow only a proportional difference.
E.3 Magnitude Tests
@@
In general, we observed that for the pairs of methods for which we found a significant partial ordering, we also found a significant difference in the magnitude of the different metrics. The test results for solution quality are presented in Table 14 ($\epsilon=1$) and 15 ($\epsilon=2$), for time offline in Table 16 ($\epsilon=1$) and 17 ($\epsilon=2$).
The remarkable things that we observed in the results of the magnitude tests, are highlighted per performance metric in the following subsections.
E.3.1 Magnitude Tests Solution Quality
In Table 14 and Table 15 a comparison of the magnitude differences in solution quality (makespan) on the double hits are provided. Here, the normalized averages indicate how much this difference is. We see that the $\mathsf{stnu}$ method for all pair-wise comparisons on the double hits results in the lowest normalized average of the makespan. In the comparison of $\mathsf{reactive}$ and $\mathsf{proactive}$ we see lower normalized averages for the $\mathsf{reactive}$ approach. Furthermore, we observe that for $\epsilon=2$ the magnitude differences become larger for the different pairs.
$[n]\ t\ (p)$ [260] -9.367 (*) [260] -13.762 (*) [260] -13.703 (*) [340] -10.114 (*) [340] -10.542 (*) [340] -5.637 (*)
norm. avg. $\mathsf{stnu}$: 0.974 $\mathsf{stnu}$: 0.962 $\mathsf{stnu}$: 0.962 $\mathsf{react}$: 0.986 $\mathsf{react}$: 0.985 $\mathsf{pro}_{\mathsf{SAA}}$: 0.999
norm. avg. $\mathsf{react}$: 1.026 $\mathsf{pro}_{\mathsf{SAA}}$: 1.038 $\mathsf{pro}_{0.9}$: 1.038 $\mathsf{pro}_{\mathsf{SAA}}$: 1.014 $\mathsf{pro}_{0.9}$: 1.015 $\mathsf{pro}_{0.9}$: 1.001
$[n]\ t\ (p)$ [230] -7.62 (*) [230] -8.608 (*) [230] -8.992 (*) [270] -6.868 (*) [270] -7.713 (*) [270] -5.82 (*)
norm. avg. $\mathsf{stnu}$: 0.982 $\mathsf{stnu}$: 0.978 $\mathsf{stnu}$: 0.977 $\mathsf{react}$: 0.995 $\mathsf{react}$: 0.993 $\mathsf{pro}_{\mathsf{SAA}}$: 0.998
norm. avg. $\mathsf{react}$: 1.018 $\mathsf{pro}_{\mathsf{SAA}}$: 1.022 $\mathsf{pro}_{0.9}$: 1.023 $\mathsf{pro}_{\mathsf{SAA}}$: 1.005 $\mathsf{pro}_{0.9}$: 1.007 $\mathsf{pro}_{0.9}$: 1.002
$[n]\ t\ (p)$ [280] -4.072 (*) [280] -5.946 (*) [280] -7.06 (*) [320] -5.698 (*) [320] -9.422 (*) [320] -6.226 (*)
norm. avg. $\mathsf{stnu}$: 0.993 $\mathsf{stnu}$: 0.99 $\mathsf{stnu}$: 0.988 $\mathsf{react}$: 0.997 $\mathsf{react}$: 0.995 $\mathsf{pro}_{\mathsf{SAA}}$: 0.998
norm. avg. $\mathsf{react}$: 1.007 $\mathsf{pro}_{\mathsf{SAA}}$: 1.01 $\mathsf{pro}_{0.9}$: 1.012 $\mathsf{pro}_{\mathsf{SAA}}$: 1.003 $\mathsf{pro}_{0.9}$: 1.005 $\mathsf{pro}_{0.9}$: 1.002
$[n]\ t\ (p)$ [330] -11.359 (*) [330] -11.391 (*) [330] -11.346 (*) [370] -2.858 (*) [370] -6.734 (*) [370] -6.014 (*)
$[n]\ t\ (p)$ [358] -6.316 (*) [370] -6.813 (*) [370] -6.817 (*) [388] -2.833 (*) [388] -7.738 (*) [400] -0.514 (0.608)
norm. avg. $\mathsf{stnu}$: 0.991 $\mathsf{stnu}$: 0.99 $\mathsf{stnu}$: 0.99 $\mathsf{react}$: 0.999 $\mathsf{react}$: 0.999 $\mathsf{pro}_{\mathsf{SAA}}$: 1.0
norm. avg. $\mathsf{react}$: 1.009 $\mathsf{pro}_{\mathsf{SAA}}$: 1.01 $\mathsf{pro}_{0.9}$: 1.01 $\mathsf{pro}_{\mathsf{SAA}}$: 1.001 $\mathsf{pro}_{0.9}$: 1.001 $\mathsf{pro}_{0.9}$: 1.0
Table 14: Magnitude test on solution quality for noise factor $\epsilon=1$. Using a pairwise t-test, including all instances for which both methods found a feasible solution. Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.
$[n]\ t\ (p)$ [205] -9.128 (*) [204] -16.872 (*) [206] -17.395 (*) [230] -10.858 (*) [233] -11.811 (*) [231] -4.856 (*)
norm. avg. $\mathsf{stnu}$: 0.963 $\mathsf{stnu}$: 0.929 $\mathsf{stnu}$: 0.925 $\mathsf{react}$: 0.966 $\mathsf{react}$: 0.964 $\mathsf{pro}_{\mathsf{SAA}}$: 0.997
norm. avg. $\mathsf{react}$: 1.037 $\mathsf{pro}_{\mathsf{SAA}}$: 1.071 $\mathsf{pro}_{0.9}$: 1.075 $\mathsf{pro}_{\mathsf{SAA}}$: 1.034 $\mathsf{pro}_{0.9}$: 1.036 $\mathsf{pro}_{0.9}$: 1.003
$[n]\ t\ (p)$ [170] -10.855 (*) [172] -11.237 (*) [170] -11.89 (*) [174] -4.95 (*) [176] -6.22 (*) [173] -3.656 (*)
norm. avg. $\mathsf{stnu}$: 0.972 $\mathsf{stnu}$: 0.968 $\mathsf{stnu}$: 0.964 $\mathsf{react}$: 0.991 $\mathsf{react}$: 0.989 $\mathsf{pro}_{\mathsf{SAA}}$: 0.997
norm. avg. $\mathsf{react}$: 1.028 $\mathsf{pro}_{\mathsf{SAA}}$: 1.032 $\mathsf{pro}_{0.9}$: 1.036 $\mathsf{pro}_{\mathsf{SAA}}$: 1.009 $\mathsf{pro}_{0.9}$: 1.011 $\mathsf{pro}_{0.9}$: 1.003
$[n]\ t\ (p)$ [140] -4.151 (*) [148] -6.017 (*) [138] -6.149 (*) [152] -4.199 (*) [160] -6.746 (*) [150] -5.404 (*)
norm. avg. $\mathsf{stnu}$: 0.99 $\mathsf{stnu}$: 0.986 $\mathsf{stnu}$: 0.985 $\mathsf{react}$: 0.994 $\mathsf{react}$: 0.991 $\mathsf{pro}_{\mathsf{SAA}}$: 0.996
norm. avg. $\mathsf{react}$: 1.01 $\mathsf{pro}_{\mathsf{SAA}}$: 1.014 $\mathsf{pro}_{0.9}$: 1.015 $\mathsf{pro}_{\mathsf{SAA}}$: 1.006 $\mathsf{pro}_{0.9}$: 1.009 $\mathsf{pro}_{0.9}$: 1.004
$[n]\ t\ (p)$ [141] -7.974 (*) [146] -8.139 (*) [148] -8.125 (*) [145] -3.581 (*) [155] -4.646 (*) [150] -2.823 (*)
norm. avg. $\mathsf{stnu}$: 0.984 $\mathsf{stnu}$: 0.981 $\mathsf{stnu}$: 0.979 $\mathsf{react}$: 0.998 $\mathsf{react}$: 0.997 $\mathsf{pro}_{\mathsf{SAA}}$: 0.999
norm. avg. $\mathsf{react}$: 1.016 $\mathsf{pro}_{\mathsf{SAA}}$: 1.019 $\mathsf{pro}_{0.9}$: 1.021 $\mathsf{pro}_{\mathsf{SAA}}$: 1.002 $\mathsf{pro}_{0.9}$: 1.003 $\mathsf{pro}_{0.9}$: 1.001
$[n]\ t\ (p)$ [94] -0.305 (0.761) [114] -0.296 (0.767) [114] -0.93 (0.355) [76] -1.573 (0.12) [99] -4.187 (*) [94] 0.59 (0.556)
norm. avg. $\mathsf{stnu}$: 0.999 $\mathsf{stnu}$: 0.999 $\mathsf{stnu}$: 0.998 $\mathsf{react}$: 0.999 $\mathsf{react}$: 0.999 $\mathsf{pro}_{\mathsf{SAA}}$: 1.0
norm. avg. $\mathsf{react}$: 1.001 $\mathsf{pro}_{\mathsf{SAA}}$: 1.001 $\mathsf{pro}_{0.9}$: 1.002 $\mathsf{pro}_{\mathsf{SAA}}$: 1.001 $\mathsf{pro}_{0.9}$: 1.001 $\mathsf{pro}_{0.9}$: 1.0
Table 15: Magnitude test on solution quality for noise factor $\epsilon=2$. Using a pairwise t-test, including all instances for which both methods found a feasible solution. Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.
E.3.2 Magnitude Tests Time Offline
@@
For time offline, $\epsilon=2$, ubo50 and ubo100, the Wilcoxon test found that the $\mathsf{stnu}$ is better than e.g. $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$, while looking at the magnitude results on double hits (this is an important difference in test procedure!), we find that the $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ are much faster offline than $\mathsf{stnu}$, which is also expected. Then, we observe no difference between $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$, which can be explained by the fact that both have the same offline procedure and we select only double hits for the test.
$[n]\ t\ (p)$ [340] nan (nan) [260] -365.854 (*) [340] -77.22 (*) [260] -365.854 (*) [340] -77.22 (*) [260] -19.399 (*)
norm. avg. $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.14 $\mathsf{react}$: 0.32 $\mathsf{pro}_{0.9}$: 0.14 $\mathsf{pro}_{0.9}$: 0.32 $\mathsf{pro}_{\mathsf{SAA}}$: 0.65
norm. avg. $\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.86 $\mathsf{pro}_{\mathsf{SAA}}$: 1.68 $\mathsf{stnu}$: 1.86 $\mathsf{pro}_{\mathsf{SAA}}$: 1.68 $\mathsf{stnu}$: 1.35
$[n]\ t\ (p)$ [270] nan (nan) [230] -45.769 (*) [270] -52.707 (*) [230] -45.769 (*) [270] -52.707 (*) [230] -3.062 (*)
norm. avg. $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.22 $\mathsf{react}$: 0.28 $\mathsf{pro}_{0.9}$: 0.22 $\mathsf{pro}_{0.9}$: 0.28 $\mathsf{pro}_{\mathsf{SAA}}$: 0.91
norm. avg. $\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.78 $\mathsf{pro}_{\mathsf{SAA}}$: 1.72 $\mathsf{stnu}$: 1.78 $\mathsf{pro}_{\mathsf{SAA}}$: 1.72 $\mathsf{stnu}$: 1.09
$[n]\ t\ (p)$ [320] nan (nan) [280] -38.703 (*) [320] -53.836 (*) [280] -38.703 (*) [320] -53.836 (*) [280] 0.5 (0.618)
norm. avg. $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.3 $\mathsf{react}$: 0.32 $\mathsf{pro}_{0.9}$: 0.3 $\mathsf{pro}_{0.9}$: 0.32 $\mathsf{stnu}$: 1.02
norm. avg. $\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.7 $\mathsf{pro}_{\mathsf{SAA}}$: 1.68 $\mathsf{stnu}$: 1.7 $\mathsf{pro}_{\mathsf{SAA}}$: 1.68 $\mathsf{pro}_{\mathsf{SAA}}$: 0.98
$[n]\ t\ (p)$ [370] nan (nan) [330] -36.15 (*) [370] -72.563 (*) [330] -36.15 (*) [370] -72.563 (*) [330] -5.636 (*)
norm. avg. $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.36 $\mathsf{react}$: 0.24 $\mathsf{pro}_{0.9}$: 0.36 $\mathsf{pro}_{0.9}$: 0.24 $\mathsf{stnu}$: 0.82
norm. avg. $\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.64 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76 $\mathsf{stnu}$: 1.64 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76 $\mathsf{pro}_{\mathsf{SAA}}$: 1.18
$[n]\ t\ (p)$ [388] nan (nan) [358] -27.209 (*) [388] -70.349 (*) [370] -25.293 (*) [400] -70.926 (*) [370] -6.391 (*)
norm. avg. $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.43 $\mathsf{react}$: 0.27 $\mathsf{pro}_{0.9}$: 0.45 $\mathsf{pro}_{0.9}$: 0.27 $\mathsf{stnu}$: 0.79
norm. avg. $\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.57 $\mathsf{pro}_{\mathsf{SAA}}$: 1.73 $\mathsf{stnu}$: 1.55 $\mathsf{pro}_{\mathsf{SAA}}$: 1.73 $\mathsf{pro}_{\mathsf{SAA}}$: 1.21
Table 16: Magnitude test on time offline for noise factor $\epsilon=1$. Using a pairwise t-test, including all instances for which both methods found a feasible solution. Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.
$[n]\ t\ (p)$ [233] nan (nan) [205] -229.689 (*) [230] -59.39 (*) [206] -230.812 (*) [231] -59.404 (*) [204] -17.066 (*)
norm. avg. $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.15 $\mathsf{react}$: 0.35 $\mathsf{pro}_{0.9}$: 0.15 $\mathsf{pro}_{0.9}$: 0.35 $\mathsf{pro}_{\mathsf{SAA}}$: 0.61
norm. avg. $\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.85 $\mathsf{pro}_{\mathsf{SAA}}$: 1.65 $\mathsf{stnu}$: 1.85 $\mathsf{pro}_{\mathsf{SAA}}$: 1.65 $\mathsf{stnu}$: 1.39
$[n]\ t\ (p)$ [176] nan (nan) [170] -82.607 (*) [174] -63.732 (*) [170] -82.607 (*) [173] -63.382 (*) [172] 0.379 (0.705)
norm. avg. $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.21 $\mathsf{react}$: 0.23 $\mathsf{pro}_{0.9}$: 0.21 $\mathsf{pro}_{0.9}$: 0.23 $\mathsf{stnu}$: 1.01
norm. avg. $\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.79 $\mathsf{pro}_{\mathsf{SAA}}$: 1.77 $\mathsf{stnu}$: 1.79 $\mathsf{pro}_{\mathsf{SAA}}$: 1.77 $\mathsf{pro}_{\mathsf{SAA}}$: 0.99
$[n]\ t\ (p)$ [160] nan (nan) [140] -33.765 (*) [152] -40.336 (*) [138] -33.253 (*) [150] -39.904 (*) [148] -2.168 (*)
norm. avg. $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.34 $\mathsf{react}$: 0.33 $\mathsf{pro}_{0.9}$: 0.34 $\mathsf{pro}_{0.9}$: 0.33 $\mathsf{stnu}$: 0.9
norm. avg. $\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.66 $\mathsf{pro}_{\mathsf{SAA}}$: 1.67 $\mathsf{stnu}$: 1.66 $\mathsf{pro}_{\mathsf{SAA}}$: 1.67 $\mathsf{pro}_{\mathsf{SAA}}$: 1.1
$[n]\ t\ (p)$ [148] 26.653 (*) [141] 27.168 (*) [146] -0.952 (0.343) [155] nan (nan) [150] -45.457 (*) [145] -44.29 (*)
norm. avg. $\mathsf{stnu}$: 1.7 $\mathsf{stnu}$: 1.72 $\mathsf{stnu}$: 0.95 $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{pro}_{0.9}$: 0.24 $\mathsf{react}$: 0.24
norm. avg. $\mathsf{pro}_{0.9}$: 0.3 $\mathsf{react}$: 0.28 $\mathsf{pro}_{\mathsf{SAA}}$: 1.05 $\mathsf{react}$: 1.0 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76
$[n]\ t\ (p)$ [114] 14.765 (*) [94] 15.183 (*) [114] -3.125 (*) [99] nan (nan) [94] -42.151 (*) [76] -42.944 (*)
norm. avg. $\mathsf{stnu}$: 1.55 $\mathsf{stnu}$: 1.61 $\mathsf{stnu}$: 0.83 $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{pro}_{0.9}$: 0.21 $\mathsf{react}$: 0.18
norm. avg. $\mathsf{pro}_{0.9}$: 0.45 $\mathsf{react}$: 0.39 $\mathsf{pro}_{\mathsf{SAA}}$: 1.17 $\mathsf{react}$: 1.0 $\mathsf{pro}_{\mathsf{SAA}}$: 1.79 $\mathsf{pro}_{\mathsf{SAA}}$: 1.82
Table 17: Magnitude test on time offline for noise factor $\epsilon=2$. Using a pairwise t-test, including all instances for which both methods found a feasible solution. Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.
E.3.3 Magnitude Tests Time Online